# Exploring Joint Observation of the CSST Shear and clustering of astrophysical gravitational wave source measurements

### v1

@@ 4.2 Constraint results
Parameter Fiducial Value
Constraints by
CSST & ECS
Constraints by
CSST & VKI+C
Constraints by
CSST & HLKI+E
@@
Compared to the cosmological constraints derived in other works, e.g. using the dark siren method (S.-J. Jin et al., 2023), the results obtained in our work are comparable to that from the future space-based gravitational wave detector networks, e.g. Taiji (D. Cyranoski, 2016), TianQin (J. Luo et al., 2016), and Laser Interferometer Space Antenna (LISA) (P. Amaro-Seoane et al., 2017).
@@
5 Summary and Conclusion
@@
In this work, we explore a multi-messenger framework cooperating the AGWS clustering measurements of the third generation GW detector networks and the CSST shear measurement to probe the Universe. This study presents a robust forecast for cosmological constraints by synergizing the CSST cosmic shear data with AGWS clustering from next-generation detector networks, namely ECS, VKI+C, and HLKI+E configurations. The generation of mock data involves computing theoretical angular power spectra for auto- and cross-correlations, incorporating GW selection functions and lensing systematics, then adding realistic noise via Cholesky decomposition of a Gaussian covariance matrix. By using the MCMC methods, the joint analysis achieves remarkable precision, notably constraining Hubble constant $H_{0}$ to $2.2\%$ and dark energy equation of state w to $5.7\%$ with the ECS network, while also providing $\sim 5\%$ accuracy on AGWS bias parameters.
Our analysis demonstrates that the joint observation can provide stringent constraints on cosmological parameters, gravitational wave source populations, and cosmic star formation history. These findings position the next-generation GW detectors as powerful tools for the LSS cosmology and compact binary physics. In future work, the advanced GW population models, non-constant bias prescriptions, and joint analysis with CMB lensing can be considered. The methodology developed here establishes a foundation for exploiting synergies between the GW detections and galaxy surveys, paving the way for 3D cosmic LSS mapping in the multi-messenger era.
@@
P.F.S. and Y.G. acknowledge the support from National Key R&D Program of China grant Nos. 2022YFF0503404 and 2020SKA0110402, and the CAS Project for Young Scientists in Basic Research (No. YSBR-092). X.L.C. acknowledges the support of the National Natural Science Foundation of China through grant Nos. 11473044 and 11973047 and the Chinese Academy of Science grants ZDKYYQ20200008, QYZDJ- SSW-SLH017, XDB 23040100, and XDA15020200. This work is also supported by science research grants from the China Manned Space Project with grant Nos. CMS-CSST-2025-A02, CMS-CSST-2021-B01, and CMS-CSST-2021-A01. The constraint results of the photo-$z$ calibration parameters ($\Delta z$ and $\sigma_{z}$), the parameter of multiplicative error $m$, and the parameters of systematical noise ($N_{\rm sys}^{gw}$, $N_{\rm add}^{gwwl}$, and $N_{\rm add}^{wl}$) from the mock data of CSST and three GW detector configurations are shown in Figures 16 and 17.

### v2

@@ IV.2 Constraint results
Parameter Fiducial Value Constraints by CSST & ECS Constraints by CSST & VKI+C Constraints by CSST & HLKI+E
@@
Our constraint on the Hubble constant ($H_{0}=69.1^{+0.92}_{-2.1}\rm{km\;s^{-1}\;Mpc^{-1}}$ , i.e., $2.2\%$ precision) represents a significant improvement over current gravitational-wave-based measurements. For example, the most recent analysis from the LIGO–Virgo–KAGRA Collaboration using O4a data (The LIGO Scientific Collaboration et al., 2025) yields $H_{0}=81.6^{+21.5}_{-15.9}\rm{km\;s^{-1}\;Mpc^{-1}}$ from the dark siren measurement alone, $H_{0}=78.4^{+25.7}_{-12.0}\rm{km\;s^{-1}\;Mpc^{-1}}$ from bright siren GW170817, and $H_{0}=76.4^{+23.0}_{-18.1}\rm{km\;s^{-1}\;Mpc^{-1}}$ from spectral siren measurement. Besides, compared to the cosmological constraints derived in other forecast works, e.g. using the dark siren method (S.-J. Jin et al., 2023), our results are comparable to that from the future space-based gravitational wave detector networks, e.g. Taiji (D. Cyranoski, 2016), TianQin (J. Luo et al., 2016), and Laser Interferometer Space Antenna (LISA) (P. Amaro-Seoane et al., 2017).
@@
To better understand the origin of our cosmological constraints, we decompose the MCMC posterior covariance by data type and redshift bin combination. We find that the cosmological parameters sensitive to geometry and structure growth (e.g., $h$, $w$, $\Omega_{m}$, $\Omega_{b}$, $\sigma_{8}$, and $n_{s}$) are mainly constrained by the shear auto-power spectrum $C^{wlwl}(\ell)$ and cross-power spectrum $C^{gwwl}(\ell)$, which are sensitive to the matter fluctuations and expansion of the Universe. The AGWS auto-power spectrum $C^{gwgw}(\ell)$ primarily constrains the parameters of GW bias ($b_{i}^{gw}$) and the cosmic star formation rate history. When combined with $C^{gwwl}$, the degeneracies between parameters can be effectively broken and more accurate constraint results can be obtained. Regarding the tomographic bin combinations, we find that the most informative cross-correlations are adjacent and near-diagonal bins (e.g., GW bin 2 $\times$ CSST bin 2–3, and GW bin 3 $\times$ CSST bin 3–4), which capture the peak overlap between the AGWS redshift distribution and the shear weighting function.
V Summary and Conclusion
@@
In this work, we explore a multi-messenger framework cooperating the AGWS clustering measurements of the third generation GW detector networks and the CSST shear measurement to probe the Universe. This study presents a robust forecast for cosmological constraints by synergizing the CSST cosmic shear data with AGWS clustering from next-generation detector networks, namely ECS, VKI+C, and HLKI+E configurations. The generation of mock data involves computing theoretical angular power spectra for auto- and cross-correlations, incorporating GW selection functions and lensing systematics, then adding realistic noise via Cholesky decomposition of a Gaussian covariance matrix. By using the MCMC methods, the joint analysis achieves remarkable precision, notably constraining Hubble constant $H_{0}$ to $2.2\%$ and dark energy equation of state $w$ to $5.7\%$ with the ECS network, while also providing $\sim 5\%$ accuracy on AGWS bias parameters. Our analysis demonstrates that the joint observation can provide stringent constraints on cosmological parameters, gravitational wave source populations, and cosmic star formation history. These findings position the next-generation GW detectors as powerful tools for the LSS cosmology and compact binary physics.
We also notice that some assumptions made in this work can be improved to obtain more reliable and accurate results, especially for the real data analysis in the future. For instance, we can adopt more realistic or measured photo-$z$ PDFs for individual sources in the CSST weak lensing survey, instead of assuming a simplified Gaussian distribution. Besides, a more reasonable GW detection rate for the BHNS systems can be used in the future to further enhance the reliability of the results. Other effects, such as GW detector sensitivity variations across the sky and galaxy survey masks, also need to be carefully considered in the analysis of real data. Despite these challenges, the methodology developed in our work establishes a foundation for exploiting synergies between the AGWS detections and weak lensing surveys, paving the way for the LSS mapping in the multi-messenger era.
@@
P.F.S. and Y.G. acknowledge the support from the CAS Project for Young Scientists in Basic Research (No. YSBR-092), and National Key R&D Program of China grant Nos. 2022YFF0503404 and 2020SKA0110402. X.L.C. acknowledges the support of the National Natural Science Foundation of China through grant Nos. 11473044 and 11973047 and the Chinese Academy of Science grants ZDKYYQ20200008, QYZDJ- SSW-SLH017, XDB 23040100, and XDA15020200. This work is also supported by science research grants from the China Manned Space Project with grant Nos. CMS-CSST-2025-A02, CMS-CSST-2021-B01, and CMS-CSST-2021-A01.
The constraint results of the photo-$z$ calibration parameters ($\Delta z$ and $\sigma_{z}$), the parameter of multiplicative error $m$, and the parameters of systematical noise ($N_{\rm sys}^{gw}$, $N_{\rm add}^{gwwl}$, and $N_{\rm add}^{wl}$) from the mock data of CSST and three GW detector configurations are shown in Figures 16 and 17.